hi, can you please help me with a homework problem, ive got a hint, but im not sure if its division, I will be thankful with a little help to verify my math h.w... well! here it goes
" A certain rectangle has a perimeter of 26cm. Its length is 7.25cm.
Write an equation, and find the width".
thanks !!

Answer by cdafonseca(2)

perimeter=w+w+h+h
26=7.25+7.25+h+h
26=14.5+2h
26-14.5=2h
14.5=2h
14.5/2=h
5.75